OpenStudy (anonymous):

Celina is buying fabric scraps to make koozies for her glasses. Each glass is cylindrical with a diameter of 3 inches and a height of 6 inches. The koozies will cover only the sides of the glasses. What is the amount of fabric she will use? What does this amount represent? Use for . Show your work.

OpenStudy (anonymous):

\[SA=2\pi r h+ \pi r^2\] is surface area with both circles out because she wont cover the bottm"only the sides" d=2r=3 r=3/2 h=6

OpenStudy (anonymous):

\[SA=2 \pi r h\]

OpenStudy (anonymous):

\[2 \pi (3/2)6\]

OpenStudy (anonymous):

thanks !